OpenStudy (anonymous):

Solve the following equation: 9x -3.2 = 6.8 + 8x

OpenStudy (anonymous):

put the like terms together 9x-8x=6.8+3.2 x=10

OpenStudy (anonymous):

9x - 3.2 = 6.8 + 8x +3.2 +3.2 9x = 10 + 8x -8x -8x x = 10